I am trying to write an equation based on table data.
X: -2, -1, 0, 1, 2
Y: -8, -2, 0, -2, -8
Based on the differences, I know that it should be a quadratic equation, but I do not know the proper format. How do you take the data from the table to create the equation?

Answer by stanbon(75887)

I am trying to write an equation based on table data.
X: -2, -1, 0, 1, 2
Y: -8, -2, 0, -2, -8
-------------------------------
Form: y = ax^2 + bx + c
Substitute three of the x/y pairs; then solve for a, b, c:
-8 = 4a -2b + c
-2 = a -b + c
0 = 0a +0b + c
-----------------------
Using the matrix function of a TI calculator I get:
a = -2 ; b = 0 ; c = 0
-------------------
Equation:
y = -2x^2
===================
Cheers,
Stan H.